You have a savings bond that $1000 when it matures in 20 years from today but you need cash now. If the current interest rate is 4% what can you get for the bond if you sell today?
Answer #1

This is like a zero coupon bond, that is not paying any intermittent interest.

Zero Coupon Bond Value = 7 (1+r) F = face value of bond r=rate or yield t=time to maturity

1000 Value = +0,04)

Value = 1000/2.1911

Value = 456.39